OpenStudy (anonymous):

Can anyone help?! suppose f(x)=4x-2 and g(x)=-2x+1 find the value of f(5)/g(-3)

OpenStudy (anonymous):

1.Plug in x=5 to f(x) to get number #1 2. Plug in x=-3 into g(x) to get number #2 Then divide number #1/number #2

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I plugged in 5 and -3 into the equation then divided both, why is it wrong?

OpenStudy (anonymous):

Plug 5 in f(x)=4x-2 what did you get?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Wait I think I messed up, is the answer 2 4/7?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} f(x)=4x-2 \\ \quad \\ g(x)=-2x+1 \end{array}\qquad \begin{array}{llll} \cfrac{f({\color{brown}{ 5}})=4({\color{brown}{ 5}})-2}{g({\color{brown}{ -3}})=-2({\color{brown}{ -3}})+1} \end{array}\implies \cfrac{?}{?}\)

OpenStudy (jdoe0001):

4*5-2 = 20-2 = 18 -2*-3+1 = +6+1 = 7

OpenStudy (anonymous):

18/7= 2.57?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Yes 2(4/7) works

OpenStudy (anonymous):

Ok, thank you both!